Russell Brown

MW 11–11:50 Office: POT741

Phone: 257-3951 Messages: 257-3336 M2–3, W2–3, F 10–11 and by appointment.

Text: S. Farlow and G. Haggard, Finite Mathematics and its Applications, 2nd edition. Custom copy for University of Kentucky In this course we will learn about linear equations, linear programming and probability. We will consider how these mathematical problems arise in applications. Students will be expected to apply algorithms to solve mathematical problems and to reformulate problems described in words as mathematical equations. Course calendar:

Dates  Monday  Wednesday  Friday
Aug 23–25  2.1, 2.2
Aug 28–Sep 1 2.2, 2.3, P1 2.3, 2.4 P2
Sep 4–8  Labor day  2.5, P3
Sep 11–15  2.6, P4  3.1

Last drop without W

Sep 18–22  Review, P5 Review  Exam 1
Sep 25–29  3.2  3.3, 3.4,P6

This week, recitation will meet in the computer lab.

Oct 2–6  3.4, 3.5, P7 3.5, P8
Oct 9–13  3.6, handout 3.6, handout, P9
Oct 16–20 Review, P10 Review Exam 2

Last withdraw

Oct 23–27 5.1 5.2

Project 1 due

Oct 30–Nov 3 5.2, 5.3, P11 5.3, 6.1, P12
Nov 6–10  6.2, P13  6.2, 6.3
Nov 13–17  6.3, P14  6.4
Nov 20–24  7.3, P15  7.4  Thanksgiving vacation

Project 2 due

Nov 27–Dec 1 Review, P16 Review  Exam 3
Dec 4–8  Review  Review

Grading: Your grade will be based on 3 hour exams (300 points), a final (150 points), supplemental problems (50 points), homework and class participation (50 points) and two written projects (50 points). Tests: The test dates are marked on the syllabus. The material from the tests will be taken from lectures, the supplemental problems and the suggested problems. If you must miss an exam, please notify Brown as soon as possible. If necessary, one makeup will be given after the exam. Supplemental problems: You will be assigned 16 supplemental problems to reinforce important ideas and prepare you for exams. You will be allowed to drop the two lowest grades. These problems will be handed out over the course of the semester. Their due dates are marked on the above calendar. No late supplemental problems will be accepted for any reason at all. However, please notify Brown if you miss an assignment. You will be excused from assignments which are missed due to illness or other unforeseen problems. Homework and class participation: Most of your time in recitation will be spent going over problems from the suggested problem list below. Your grade will be based on your attendance in recitation and the number of problems you present. Projects: You will be asked to carry out two longer written assignments. Detailed descriptions of these assignments will be given approximately 4 weeks in advance of the due date. The due dates for these are indicated on the schedule above. You must work in groups of between two and five for these assignments. Suggested problems: You should work each of the following problems. You will discuss the solutions in recitation (see the above section on homework and class participation).

Section Problems
2.2  3, 13, 14, 19, 20, 26, 27, 30, 33, 34, 39
2.3  5, 7, 8, 9, 14, 15, 23, 25, 27, 29
2.4  1, 8, 10–12, 23, 25, 27, 50, 54
2.5  3, 4, 5, 10–15, 22, 23, 31, 32
2.6  1–4, 17, 24, 35–38
3.1  1, 2, 13–14, 25–28
3.2  1, 2, 7–10, 17–20, 26, 30, 34, 36
3.3  1, 2, 3, 5, 7, 9, 12
3.4  6—14, 17, 21, 26, 27, 34
3.5  1–12, 13, 20, 21, 24
3.6  22, 23, 29–35
Handout all
5.1  9–40
5.2  3–6, 16–19, 23, 24, 34–36
5.3  1–10, 27–29, 47–53, 60, 61, 65, 66
6.1  1–5, 15–17, 21–30
6.2  1–8, 15–18, 31–34, 48–52
6.3  1–12, 15, 16, 36–39, 52–53
6.4  14–17, 25–28, 41, 42
7.3  15–18, 21–31, 33, 34
7.4  7–9, 14–17, 20–23, 54–56